OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y =1/4x + 2. A. y = 4x - 1 B. y = 4x - 3 C. y = 1/4x - 3 D. y = 1/4x - 1

OpenStudy (igreen):

Parallel lines have the same slope, so what's the slope of the new line gonna be?

OpenStudy (anonymous):

@jojo4eva @felavin @Nnesha

OpenStudy (igreen):

Do you know what the slope of y = 1/4x + 2 is?

OpenStudy (anonymous):

1/4 @iGreen

OpenStudy (igreen):

Yep, and since parallel lines have THE SAME slope, what's the slope of the other line going to be?

OpenStudy (anonymous):

it would be D

OpenStudy (igreen):

Wait..it's either C or D.. Let's see. We have our point (-4, -2) and our slope 1/4. Plug them into the Point-Slope Form: \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value, \(x_1\) is the x-value, and \(m\) is the slope. \(y +2 = \dfrac{1}{4} (x +4)\) Distribute \(\dfrac{1}{4}\) into the parenthesis. \(y + 2 = \dfrac{1}{4} \! \! x + 1\) Subtract 2 to both sides, what's 1 - 2? @IndianBoyNiko

OpenStudy (anonymous):

1 , it's D

OpenStudy (igreen):

You mean -1? That gives us \(y = \dfrac{1}{4} \! \! x - 1\). So you're correct, your answer is D..

OpenStudy (anonymous):

thanks